OpenStudy (shamil98):

The function h = -t2 + 95 models the path of a ball thrown by a boy where h represents height, in feet, and t represents the time, in seconds, that the ball is in the air. Assuming the boy lives at sea level where h = 0 ft, which is a likely place the boy could have been standing when he threw this ball? A. his backyard B. an underground cave C. a ladder D. a bridge

OpenStudy (shamil98):

@wio

OpenStudy (shamil98):

Take a trip down memory lane :D

OpenStudy (anonymous):

Let \(t=0\) to see where the ball is when he threw it.

OpenStudy (shamil98):

The given is h = -t^2 + 95 he lives at sea level so h = 0

OpenStudy (anonymous):

No, I'm saying find \(h(0\). Since \(t=0\) when he threw it.

OpenStudy (shamil98):

The ball is in the air for t seconds as well.

OpenStudy (anonymous):

Baiscaly \[ h(0) = (0)^2+95=95 \]

OpenStudy (anonymous):

So when he threw the ball he was 95 ft above sea level.

OpenStudy (shamil98):

Mhm.

OpenStudy (anonymous):

So a ladder likely... or a really high up bridge

OpenStudy (shamil98):

I would say a bridge.. or a ladder, but i doubt he would be on a ladder that's 95 ft

OpenStudy (anonymous):

Whatever, who cares.